# PPT: Token Pruning and Pooling for Efficient Vision Transformers

### v1

@@ Abstract
Vision Transformers (ViTs) have emerged as powerful models in the field of computer vision, delivering superior performance across various vision tasks. However, the high computational complexity poses a significant barrier to their practical applications in real-world scenarios. Motivated by the fact that not all tokens contribute equally to the final predictions and fewer tokens bring less computational cost, reducing redundant tokens has become a prevailing paradigm for accelerating vision transformers. However, we argue that it is not optimal to either only reduce inattentive redundancy by token pruning, or only reduce duplicative redundancy by token merging. To this end, in this paper we propose a novel acceleration framework, namely token Pruning & Pooling Transformers (PPT), to adaptively tackle these two types of redundancy in different layers. By heuristically integrating both token pruning and token pooling techniques in ViTs without additional trainable parameters, PPT effectively reduces the model complexity while maintaining its predictive accuracy. For example, PPT reduces over 37% FLOPs and improves the throughput by over 45% for DeiT-S without any accuracy drop on the ImageNet dataset.
@@ 1 Introduction
In recent years, vision transformers (ViTs) have demonstrated promising results in many vision tasks such as image classification (Dosovitskiy et al., 2021; Jiang et al., 2021; Touvron et al., 2021; Yuan et al., 2021), object detection (Carion et al., 2020; Dai et al., 2021; Li et al., 2022; Zhu et al., 2021), and semantic segmentation (Kirillov et al., 2023; Liu et al., 2021; Wang et al., 2021). Compared with the convolutional neural networks (CNNs), ViTs have the property of modeling long-range dependencies with the attention mechanism (Vaswani et al., 2017), which introduces fewer inductive biases hence has the potential to absorb more training data. However, densely modeling long-range dependencies among image tokens can lead to computational inefficiency, especially when dealing with large datasets and training iterations (Carion et al., 2020; Dosovitskiy et al., 2021). This in turn limits further implementation of ViTs in real-world scenarios.
Given the strong correlation between model complexity and the number of tokens in ViTs, a direct approach to accelerate ViTs is to reduce the number of redundant tokens. Moreover, some studies also showed that not all tokens contribute equally to the final predictions (Caron et al., 2021; Pan et al., 2021) . Existing attempts to achieve token compression mainly consist of two branches of solutions, i.e., token pruning and token pooling. The former approach emphasizes the design of different importance evaluation strategies to identify and retain relevant tokens while discarding irrelevant ones, as demonstrated in previous research (Fayyaz et al., 2022; Liang et al., 2022; Rao et al., 2021; Xu et al., 2022; Tang et al., 2022). The latter technique primarily concentrates on merging similar image tokens using a predefined similarity evaluation metric and merge policy (Bolya et al., 2023; Marin et al., 2021). To summarize, there are two types of redundancy in vision transformers, i.e., inattentive redundancy and duplicative redundancy. However, it should be noted that the aforementioned each method only addresses one type of redundancy, i.e., token pruning for inattentive redundancy while token pooling for duplicative redundancy. We argue that reducing only one type of redundancy leads to suboptimal acceleration performance.
@@ 2 Related Work
Vision Transformers. Inspired by the success of Transformers in natural language processing area (Brown et al., 2020; Kenton & Toutanova, 2019; Vaswani et al., 2017), the recent advance ViT (Dosovitskiy et al., 2021) shows that the transformer can also achieve promising results in the computer vision field. However, ViT requires large-scale datasets such as ImageNet-22K and JFT-300M for model pretraining and huge computation resources. Later, DeiT (Dosovitskiy et al., 2021) addressed this issue by optimizing the training strategy and introducing a distilled token, which is designed to learn knowledge from the teacher network and improve model performance. Some following works (Chen et al., 2021; Han et al., 2021; Liu et al., 2021; Wang et al., 2021; Wu et al., 2021; Yuan et al., 2021) focus on modifying patch embedding or transformer blocks to introduce local dependencies into ViTs, resulting in significant performance improvements. LV-ViT (Jiang et al., 2021) further improves ViTs by utilizing the local information embedded in patch tokens with a new training objective called token labeling. Although the ViT and its follow-ups have achieved excellent performance and demonstrated their strong potential as an alternative to CNNs, the high computational costs remain a challenge for practical implementation in real-world scenarios. The computational cost of ViTs is mainly attributed to the quadratic and linear time complexity of the multi-head self-attention (MSA) and feed-forward network (FFN), respectively, with respect to the number of tokens. As a result, researchers have been exploring token compression as a prevalent paradigm for accelerating ViTs, which can effectively alleviate the computational burden while retaining their expressive power and performance.
@@
Token Pooling. On the other hand, there are some attempts that recognize the significance of pooling tokens together. To alleviate information loss, Evo-ViT (Xu et al., 2022) and EViT (Liang et al., 2022) merge the tokens they pruned into a single token. To improve the efficiency of ViTs, Token Pooling (Marin et al., 2021) utilizes a K-Means-based clustering approach to exploit redundancies in the images, while ToMe (Bolya et al., 2023) thoroughly investigates token similarity and proposes a Bipartite Soft Matching algorithm (BSM) to gradually merge similar tokens. Token pooling methods can reduce the duplicative redundancy, but they do not take into account that not all tokens contribute equally to the final prediction.
@@ 3.2 Token Pruning for Inattentive Redundancy
Token Scoring. There are quite a few non-parametric scoring mechanisms for tokens in prior work, e.g., the attention scores $\bm{A}_{cls}$ between the classification token and image tokens (Liang et al., 2022; Xu et al., 2022). Furthermore, ATS (Fayyaz et al., 2022) takes the values matrix $\bm{V}$ into consideration and proposes a more comprehensive token scoring metric:
${\bm{A}}=Softmax(\frac{\bm{Q}\cdot\bm{K}^{T}}{\sqrt{d}}),\quad\mbox{Score}_{i}=\frac{{\bm{A}_{1,i+1}\times\left\lVert\bm{V}_{i+1}\right\rVert}}{\sum_{j=1}^{N}\bm{A}_{1,j+1}\times\left\lVert\bm{V}_{j+1}\right\rVert}.$ (1)
Token Selecting. For token selection, we return to the traditional Top-K selection policy for more controllable compression ratio, i.e., preserve the Top-K important tokens while remove the other inattentive tokens, which is widely used in many works (Liang et al., 2022; Rao et al., 2021).
@@ 3.3 Token Pooling for Duplicative Redundancy
The token pooling techniques aims to merge similar image tokens together and can decrease the duplicative redundancy in the model. Recently, the Bipartite Soft Matching algorithm (BSM) algorithm (Bolya et al., 2023) shows superior performance in token merging. BSM first partitions the tokens into two sets of roughly equal size. It then draws an edge from each token in one set to the token in the other set with the highest $cosine$ similarity score. The top-K-similar edges are selected, and tokens that are still connected are merged by averaging their features. In addition, it is necessary to maintain a variable $\bm{s}$ that tracks the size of the tokens in order to minimize information loss. $\bm{s}$ is a row vector that reflects the number of tokens represented and are combined with tokens any time. In addition, it is also used as a weight to reflect its importance in the calculation of attention matrix:
@@
We also explore another clustering-based token pooling method in Section C of appendix and demonstrate the BSM performer better in our framework.
@@ 3.4 Token Pruning & Pooling Transformer
A Closer Look at Token Pruning and Token Pooling. We first conduct some analysis of the importance scores in different layers. Our research reveals an intriguing phenomenon: the variance of significance scores assigned to image tokens for each sample increases as the number of layers in the model increases, as depicted in Figure 3 (deeper analysis are shown in Section B of the appendix). This suggests that the importance of image tokens becomes more distinct as the layer deepens. As a result, token pruning techniques may be preferred at deeper layers, where certain tokens exhibit significantly lower importance scores and are therefore more likely to be pruned. Meanwhile, premature token pruning at shallow layers may result in irreversible information loss and negatively impact model performance.
@@
${S_{op}}_{i}=var(\mbox{Score}_{i}),\quad\mbox{Policy}_{i}=\begin{cases}\mbox{Token Pruning},\mbox{if }{S_{op}}_{i}>\tau\\ \mbox{Token Pooling},\mbox{otherwise}\end{cases},$ (3)
4 Experiments
In this section, we empirically investigate the superiority of the proposed PPT through extensive experiments on ImageNet-1K (ILSVRC2012) (Deng et al., 2009), which contains approximately 1.28M training images and 50K validation images. We compare our proposed model with state-of-the-art models and conduct thorough ablation studies to gain a better understanding of its effectiveness.
[…]
4.1 Main Results
Table 1: Comparison of the accelerated vision transformers with different methods applied to multiple vanilla ViTs on ImageNet. The model with ${}^{\dagger}$’ is trained from scratch (300 epochs).
@@ 4.1 Main Results
Comparisons with existing token compression methods. Even though we do not add extra parameters as Rao et al. (2021), or apply complicated token reorganization tricks as Liang et al. (2022); Xu et al. (2022), the experimental results, as presented in Table 1 and Figure 4, demonstrate that our method achieves higher accuracy with comparable computation cost. Specifically, PPT reduces FLOPs by over 37% and improves throughput by over 45% without any accuracy drop on the classic DeiT-S model as shown in Table 1. Furthermore, the superiority of PPT is evident across various FLOPs when compared with other token compression methods, as shown in Figure 4. Specifically, our method excels in the following three aspects: Firstly, at lower FLOPs (below 2.5G), our method showcases a noteworthy performance improvement of 0.7%-3.8% compared to other methods. Secondly, at higher FLOPs (3.0G above), our method outperforms the baseline model and surpasses the capabilities of the comparison methods. Lastly, when used in an off-the-shelf setting, PPT demonstrates significant improvement of 0.1%-2.3% across various FLOPs compared to other methods, achieving competitive results with those obtained by fine-tuned models.
@@
Figure 4: Comparison between our method and other methods under different FLOPs. We conducted a comprehensive comparison of the performance of various methods after fine-tuning (left) and off-the-shelf (right), which highlights the superior performance of our method.
Application on different variants of ViTs. Apart from the standard ViTs (Dosovitskiy et al., 2021; Touvron et al., 2021), subsequent studies (Chen et al., 2021; Han et al., 2021; Heo et al., 2021; Jiang et al., 2021; Liu et al., 2021; Radosavovic et al., 2020; Wang et al., 2021; Wu et al., 2021; Xu et al., 2021; Yuan et al., 2021; Yue et al., 2021) further improve the performance of ViTs by varying the initial architecture or optimization strategies. To further showcase the potential and superiority of our approach, we extend the integration of PPT with the LV-ViT-S. As shown in Table 2 and Figure 5, our PPT-LV-S can achieve a highly competitive performance among numerous vision transformers from the perspective of accuracy-computation trade-off, and further demonstrate superiority over other token compression methods on LV-ViT-S. To further demonstrate the generality and superiority of our approach, we also apply PPT on more variants of ViTs, such as T2T-ViT and PS-ViT, as shown in Figure 6. More detailed results are shown in Section B of the appendix.
@@
Table 2: Comparisons with different variants of ViTs on ImageNet. We compress the LV-ViT-S (Jiang et al., 2021) as the base model and achieve promising accuracy-FLOPs trade-off. Model Params FLOPs Top-1 (M) (G) Acc. (%) DeiT-S 22.1 4.6 79.8 DeiT-B 86.6 17.6 81.8 PVT-Small 24.5 3.8 79.8 PVT-Medium 44.2 6.7 81.2 CoaT-Lite Small 20.0 4.0 81.9 CrossViT-S 26.7 5.6 81.0 Swin-T 29.0 4.5 81.3 Swin-S 50.0 8.7 83.0 T2T-ViT-14 22.0 4.8 81.5 T2T-ViT-24 64.1 14.1 82.3 RegNetY-8G 39.0 8.0 81.7 RegNetY-16G 84.0 16.0 82.9 PS-ViT-B/14 21.3 5.4 81.5 PS-ViT-B/18 21.3 8.8 82.3 PiT-S 23.5 2.9 80.9 PiT-B 73.8 12.5 82.0 CvT-13 20.0 4.5 81.6 TNT-S 23.8 5.2 81.5 TNT-B 66.0 14.1 82.9 LV-ViT-S 26.2 6.6 83.3 DynamicViT-LV-S 26.9 4.6 83.0 PS-LV-ViT-S 26.2 4.7 82.4 EViT-LV-S 26.2 4.7 83.0 PPT-LV-S (Ours) (off-the-shelf) 25.8 4.6 82.8 PPT-LV-S (Ours) 25.8 4.6 83.1
@@
[Uncaptioned image] Figure 5: Comparison of different models with various accuracy-FLOPs trade-off. Our PPT-LV-S achieves a quite competitive trade-off than other ViTs.
Figure 6: The performance of applying our method to more ViTs (without fine-tuning).
@@
Figure 7: Visualizations of token compression results on DeiT-S. The masked regions represent the inattentive redundancy and are pruned, while the patches with the same inner and border color means the duplicative redundancy and are pooled. The redundancy in the image is pyramid reduced, and our method can adaptively execute different strategies at different stages based on the input. We demonstrate the generalization of the model to images of varying complexity. More results are visualized in the Section D of the appendix.
@@ 4.2 Ablation Study
Random Rule-Based Adaptive (Ours) Policy Inversion
@@
We also explore the impact of different pruning and pooling policy and different metrics for redundancy discrimination applied in our framework, we show the results in Section C due to the space limitations.

### v2

@@ Abstract
Vision Transformers (ViTs) have emerged as powerful models in the field of computer vision, delivering superior performance across various vision tasks. However, the high computational complexity poses a significant barrier to their practical applications in real-world scenarios. Motivated by the fact that not all tokens contribute equally to the final predictions and fewer tokens bring less computational cost, reducing redundant tokens has become a prevailing paradigm for accelerating vision transformers. However, we argue that it is not optimal to either only reduce inattentive redundancy by token pruning, or only reduce duplicative redundancy by token merging. To this end, in this paper we propose a novel acceleration framework, namely token Pruning & Pooling Transformers (PPT), to adaptively tackle these two types of redundancy in different layers. By heuristically integrating both token pruning and token pooling techniques in ViTs without additional trainable parameters, PPT effectively reduces the model complexity while maintaining its predictive accuracy. For example, PPT reduces over 37% FLOPs and improves the throughput by over 45% for DeiT-S without any accuracy drop on the ImageNet dataset.  [^1] The code is available at https://github.com/xjwu1024/PPT and https://github.com/mindspore-lab/models/
@@ 1 Introduction
In recent years, vision transformers (ViTs) have demonstrated promising results in many vision tasks such as image classification [8, 12, 24, 31], object detection [3, 6, 15, 33], and semantic segmentation [14, 17, 26]. Compared with the convolutional neural networks (CNNs), ViTs have the property of modeling long-range dependencies with the attention mechanism [25], which introduces fewer inductive biases hence has the potential to absorb more training data. However, densely modeling long-range dependencies among image tokens can lead to computational inefficiency, especially when dealing with large datasets and training iterations [3, 8]. This in turn limits further implementation of ViTs in real-world scenarios.
Given the strong correlation between model complexity and the number of tokens in ViTs, a direct approach to accelerate ViTs is to reduce the number of redundant tokens. Moreover, some studies also showed that not all tokens contribute equally to the final predictions [4, 19] . Existing attempts to achieve token compression mainly consist of two branches of solutions, i.e., token pruning and token pooling. The former approach emphasizes the design of different importance evaluation strategies to identify and retain relevant tokens while discarding irrelevant ones, as demonstrated in previous research [9, 16, 21, 29, 23]. The latter technique primarily concentrates on merging similar image tokens using a predefined similarity evaluation metric and merge policy [1, 18]. To summarize, there are two types of redundancy in vision transformers, i.e., inattentive redundancy and duplicative redundancy. However, it should be noted that the aforementioned each method only addresses one type of redundancy, i.e., token pruning for inattentive redundancy while token pooling for duplicative redundancy. We argue that reducing only one type of redundancy leads to suboptimal acceleration performance.
@@ 2 Related Work
Vision Transformers. Inspired by the success of Transformers in natural language processing area [2, 13, 25], the recent advance ViT [8] shows that the transformer can also achieve promising results in the computer vision field. However, ViT requires large-scale datasets such as ImageNet-22K and JFT-300M for model pretraining and huge computation resources. Later, DeiT [8] addressed this issue by optimizing the training strategy and introducing a distilled token, which is designed to learn knowledge from the teacher network and improve model performance. Some following works [5, 10, 17, 26, 27, 31] focus on modifying patch embedding or transformer blocks to introduce local dependencies into ViTs, resulting in significant performance improvements. LV-ViT [12] further improves ViTs by utilizing the local information embedded in patch tokens with a new training objective called token labeling. Although the ViT and its follow-ups have achieved excellent performance and demonstrated their strong potential as an alternative to CNNs, the high computational costs remain a challenge for practical implementation in real-world scenarios. The computational cost of ViTs is mainly attributed to the quadratic and linear time complexity of the multi-head self-attention (MSA) and feed-forward network (FFN), respectively, with respect to the number of tokens. As a result, researchers have been exploring token compression as a prevalent paradigm for accelerating ViTs, which can effectively alleviate the computational burden while retaining their expressive power and performance.
@@
Token Pooling. On the other hand, there are some attempts that recognize the significance of pooling tokens together. To alleviate information loss, Evo-ViT [29] and EViT [16] merge the tokens they pruned into a single token. To improve the efficiency of ViTs, Token Pooling [18] utilizes a K-Means-based clustering approach to exploit redundancies in the images, while ToMe [1] thoroughly investigates token similarity and proposes a Bipartite Soft Matching algorithm (BSM) to gradually merge similar tokens. Token pooling methods can reduce the duplicative redundancy, but they do not take into account that not all tokens contribute equally to the final prediction.
@@ 3.2 Token Pruning for Inattentive Redundancy
Token Scoring. There are quite a few non-parametric scoring mechanisms for tokens in prior work, e.g., the attention scores $\bm{A}_{cls}$ between the classification token and image tokens [16, 29]. Furthermore, ATS [9] takes the values matrix $\bm{V}$ into consideration and proposes a more comprehensive token scoring metric:
$\small{\bm{A}}=Softmax(\frac{\bm{Q}\cdot\bm{K}^{T}}{\sqrt{d}}),\quad\mbox{Score}_{i}=\frac{{\bm{A}_{1,i+1}\times\left\lVert\bm{V}_{i+1}\right\rVert}}{\sum_{j=1}^{N}\bm{A}_{1,j+1}\times\left\lVert\bm{V}_{j+1}\right\rVert}.$ (1)
Token Selecting. For token selection, we return to the traditional Top-K selection policy for more controllable compression ratio, i.e., preserve the Top-K important tokens while remove the other inattentive tokens, which is widely used in many works [16, 21].
@@ 3.3 Token Pooling for Duplicative Redundancy
The token pooling techniques aims to merge similar image tokens together and can decrease the duplicative redundancy in the model. Recently, the Bipartite Soft Matching algorithm (BSM) algorithm [1] shows superior performance in token merging. BSM first partitions the tokens into two sets of roughly equal size. It then draws an edge from each token in one set to the token in the other set with the highest $cosine$ similarity score. The top-K-similar edges are selected, and tokens that are still connected are merged by averaging their features.
In addition, it is necessary to maintain a variable $\bm{s}$ that tracks the size of the tokens in order to minimize information loss. $\bm{s}$ is a row vector that reflects the number of tokens represented and are combined with tokens any time. In addition, it is also used as a weight to reflect its importance in the calculation of attention matrix:
@@ 3.4 Token Pruning & Pooling Transformer
A Closer Look at Token Pruning and Token Pooling. We first conduct some analysis of the importance scores in different layers. Our research reveals an intriguing phenomenon: the variance of significance scores assigned to image tokens for each sample increases as the number of layers in the model increases, as depicted in Figure 3 This suggests that the importance of image tokens becomes more distinct as the layer deepens. As a result, token pruning techniques may be preferred at deeper layers, where certain tokens exhibit significantly lower importance scores and are therefore more likely to be pruned. Meanwhile, premature token pruning at shallow layers may result in irreversible information loss and negatively impact model performance.
@@
Table 1: Comparison of the accelerated vision transformers with different methods applied to multiple vanilla ViTs on ImageNet. The model with ${}^{\dagger}$’ is trained from scratch with 300 epochs.
@@
$\small{S_{op}}_{i}=var(\mbox{Score}_{i}),\ \mbox{Policy}_{i}=\begin{cases}\mbox{Token Pruning},\mbox{if }{S_{op}}_{i}>\tau\\ \mbox{Token Pooling},\mbox{otherwise}\end{cases},$ (3)
4 Experiments
In this section, we empirically investigate the superiority of the proposed PPT through extensive experiments on ImageNet-1K (ILSVRC2012) [7], which contains approximately 1.28M training images and 50K validation images. We compare our proposed model with state-of-the-art models and conduct thorough ablation studies to gain a better understanding of its effectiveness.
[…]
@@ 4 Experiments
Figure 4: Comparison between our method and other methods under different FLOPs. We conducted a comprehensive comparison of the performance of various methods after fine-tuning (left) and off-the-shelf (right), which highlights the superiority of our method.
@@
Table 2: Comparisons with different variants of ViTs on ImageNet. We compress the LV-ViT-S [12] as the base model and achieve promising accuracy-FLOPs trade-off. Model Params FLOPs Top-1 (M) (G) Acc. (%) DeiT-S [24] 22.1 4.6 79.8 DeiT-B [24] 86.6 17.6 81.8 PVT-Small [26] 24.5 3.8 79.8 PVT-Medium [26] 44.2 6.7 81.2 CoaT-Lite Small [28] 20.0 4.0 81.9 CrossViT-S [5] 26.7 5.6 81.0 Swin-T [17] 29.0 4.5 81.3 Swin-S [17] 50.0 8.7 83.0 T2T-ViT-14 [31] 22.0 4.8 81.5 T2T-ViT-24 [31] 64.1 14.1 82.3 RegNetY-8G [20] 39.0 8.0 81.7 RegNetY-16G [20] 84.0 16.0 82.9 PS-ViT-B/14 [32] 21.3 5.4 81.5 PS-ViT-B/18 [32] 21.3 8.8 82.3 PiT-S [11] 23.5 2.9 80.9 PiT-B [11] 73.8 12.5 82.0 CvT-13 [27] 20.0 4.5 81.6 TNT-S [10] 23.8 5.2 81.5 TNT-B [10] 66.0 14.1 82.9 LV-ViT-S [12] 26.2 6.6 83.3 DynamicViT-LV-S [21] 26.9 4.6 83.0 PS-LV-ViT-S [23] 26.2 4.7 82.4 EViT-LV-S [16] 26.2 4.7 83.0 PPT-LV-S (Ours) (off-the-shelf) 25.8 4.6 82.8 PPT-LV-S (Ours) 25.8 4.6 83.1
@@
[Uncaptioned image] Figure 5: Comparison of different models with various accuracy-FLOPs trade-off. Our PPT-LV-S achieves a quite competitive trade-off than other ViTs. [Uncaptioned image] Figure 6: The performance of applying our method to more ViTs (without fine-tuning).
@@
Figure 7: Visualizations of token compression results on DeiT-S. The masked regions represent the inattentive redundancy and are pruned, while the patches with the same inner and border color means the duplicative redundancy and are pooled. The redundancy in the image is pyramid reduced, and our method can adaptively execute different strategies at different stages based on the input. We demonstrate the generalization of the model to images of varying complexity.
4.1 Main Results
Comparisons with existing methods. Even though we do not add extra parameters as Rao et al. [21], or apply complicated token reorganization tricks as Liang et al. [16], Xu et al. [29], the experimental results, as presented in Table 1 and Figure 4, demonstrate that our method achieves higher accuracy with comparable computation cost. Specifically, PPT reduces FLOPs by over 37% and improves throughput by over 45% without any accuracy drop on the classic DeiT-S model as shown in Table 1. Furthermore, the superiority of PPT is evident across various FLOPs when compared with other token compression methods, as shown in Figure 4. Specifically, our method excels in the following three aspects: Firstly, at lower FLOPs (below 2.5 G), our method showcases a noteworthy performance improvement of 0.7%-3.8% compared to other methods. Secondly, at higher FLOPs (3.0 G above), our method outperforms the baseline model and surpasses the capabilities of the comparison methods. Lastly, when used in an off-the-shelf setting, PPT demonstrates significant improvement of 0.1%-2.3% across various FLOPs compared to other methods, achieving competitive results with those obtained by fine-tuned models.
Application on different variants of ViTs. Apart from the standard ViTs [8, 24], subsequent studies [5, 10, 11, 12, 17, 20, 26, 27, 28, 31, 32] further improve the performance of ViTs by varying the initial architecture or optimization strategies. To further showcase the potential and superiority of our approach, we extend the integration of PPT with the LV-ViT-S. As shown in Table 2 and Figure 5, our PPT-LV-S can achieve a highly competitive performance among numerous vision transformers from the perspective of accuracy-computation trade-off, and further demonstrate superiority over other token compression methods on LV-ViT-S. To further demonstrate the generality and superiority of our approach, we also apply PPT on more variants of ViTs, such as T2T-ViT and PS-ViT, as shown in Figure 6.
@@ 4.2 Ablation Study
Random Rule-Based Adaptive Policy
(Ours) Inversion